A substance believed to be an anabolic steroid (used by some athletes to build muscle) is seized from

a professional athlete and analyzed. Combustion analysis reveals that the compound is 80.0 % carbon,

9.41 % hydrogen, and 10.6 % oxygen. A mass spectrometer analysis shows that the compound has a

molar mass of 300.48 g/mol. K/U T/I C

(a) Determine the molecular formula of the compound.

(b) The substance is believed to be dianabol, a common performance-enhancing steroid. Is this analysis

data sufficient to accuse the athlete of using dianabol? Why or why not?


1
Expert's answer
2020-08-19T04:30:29-0400

as given carbon is 80%

molar mass of carbon in compound = 300*80/100 = 240 grams

no. of carbon atoms = 240/12 = 20

hydrogen is 9.41%

molar mass of Hydrogen in compound = 300*9.41/100 = 28.2 grams

no. of hydrogen atoms = 28.2/1 rougly equal to 28

and oxygen is 10.6%

molar mass of oxygen in compound is 300* 10.6/100 = 31.8

no. of oxygen atom = 31.8/16 roughly equal to 2


so molecular formula of compound is "C_{20}H_{28}O_2"


(b) No, this is not sufficient data to accuse. We need to perform chemical test in order to verify it is dianabol.